OpenStudy (anonymous):

Clarissa works in sales for a software company. Her pay each month varies since each month she receives base pay of $2,400.00 plus 8% commission on her sales. During the month of November, Clarissa had $18,000.00 in sales. For the month of November, ----- % of Clarissa's pay was from her commission.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

30%

OpenStudy (anonymous):

@hl147

OpenStudy (texaschic101):

she had 18,000 in sales....if she gets paid a commission of 8%....then she would get paid : (0.08)(2400) = 192....then add that to her base pay 2,400 + 192 = 2592 so her commission is 192 and total pay is $ 2592 192 is x% of 2592 192 = (2592)x% 192/2592 = x% 0.074 = x% 7.4 % = x I am not sure if I did this correctly, but I got 7.4%

OpenStudy (anonymous):

37.5

OpenStudy (texaschic101):

@dtan5457 ...can you take a look at this please

OpenStudy (anonymous):

Aerial's grandmother gave her $5,500.00 to save for her college education. She went to the bank to open a savings account. The bank told her they had two options available. Account A will pay 5.5% simple interest until the account is closed. Account B will pay 4.5% simple interest and if the account is left open for longer than 3 years, then at the end of the third year a bonus account will be opened with $250 that also earns 4.5% for the remainder of the time the initial account is open. When the initial account is closed, the bonus account will be closed as well and the money from the two accounts will be combined. If Aerial is planning on leaving the money in the account for 4 years and then withdrawing all funds, then account -------- is the better choice earning her ------- more.

OpenStudy (texaschic101):

I = PRT I = (5000)(0.055)(4) I = 1100 5000 + 1100 = 6100 --- this is how much she would earn for option A I = PRT I = (5000)(0.045)(4) I =900 5000 + 900 = 5900 I = PRT I = (250)(0.045)(1) I = 11.25 250 + 11.25 = 261.25 5900 + 261.25 = 6161.25 -- this is how much she would earn for option B so option A is 6100 and option B is 6161.25....better deal is option B getting her (6161.25 - 6100 = 61.25).....giving her $61.25 more.